# Simple Network Mechanism Leads to Quasi-Real Brain Activation Patterns with Drosophila Connectome

### v1

@@ 3.1 Drosophila Connectome Data Set
These statistic results are a valuable supplement to previous findings[45] and are crucial for a deeper understanding of this data set.
3.2 Model
In this section, we introduce the large scale network communication model which consumes less computational resource and can efficiently utilize the network structure. First we introduce the symbol systems in this article.
The original network presentation of neural systems often uses nodes to represent the neurons and using edges to represent the links (chemical synapses, electrical synapses, etc) between neurons[4]. However, this approach has a number of limitations: Firstly, there are more than one link (synapse) between two neurons[46] and the links are spatially and temporally asymmetric. To illustrate, two neurons shown in Fig. 3 have multiple synapses between them, and the dynamical properties of these different synapses may vary. Additionally, due to their spatial locations, the signal transmission delay and the position of action on the downstream neurons are also different, making it difficult to simply merge them.
Figure 3: An example of Complex Connection
For this reason, similar to previous research[47], we use the dual method on Drosophila connectome data set. The method is introduced below.
A dual graph is a concept in graph theory which describes a special relationship between graphs. In a dual graph, the nodes and edges of the original graph are interchanged, resulting in a new graph. In simple terms, a dual graph is obtained by replacing the nodes of the original graph with edges, and the edges with nodes. Fig. 4 is an example of an original graph and a dual graph. After being processed, the new graph can be considered to be a hyper graph, whose nodes represent the synapses and links represent the cell body. Then we can use the hyper communication model to describe it.
(a) Original Graph
(b) Dual Graph
Figure 4: An example of original graph and dual graph
In order to standardize the symbol system, we use a set to represent the neuronal cell body and use an element to represent a physical link between neurons. We regard the process of messages interactions as the elements’ interactions on sets.
Let $N=\{n_{1},...,n_{i},...\}$ represent the set of all elements and $E=\{e_{1},...,e_{j}...\}$ represent the probable sets of elements. Note that the intersection of sets include some elements, which is formalized as the following manner.
$e_{k}\cap e_{q}=\{n_{i}|(n_{i}\in e_{k})\wedge(n_{i}\in e_{q})\}$ (13)
Note that union of every element in $N$ and $E$ is the whole graph. Let $G$ represent the graph. The graph’s definition is given as follows.
$G=\bigcup_{j}e_{j}=\bigcup_{i}\{n_{i}\}.$ (14)
Because neural link is directed, let ${e_{j}}^{+}$ ( ${e_{j}}^{-}$) represent the set of elements which output (input) the signals, note that ${e_{j}}^{+}$ = ${e_{j}}^{+}\cup{e_{j}}^{-}$.
Every neuron and synapse has a activate state (state = 1) or resting state (state = 0). So set the statement of the element $i$ as $X_{i}$. $X_{i}=1$ if the element $n_{i}$ is active, otherwise $X_{i}=0$. Then let $Y_{j}$ represent the statement of set $e_{j}$. $Y_{j}=1$ if the set $e_{j}$ is active, otherwise $Y_{j}=0$.
We use a linear threshold activation function. Then, general transmission function is defined in the following manner,
$Y_{j}=\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})$ (15)
where $\omega(n_{k},e_{i},e_{j})$ represents the link weight between set $i$ and set $j$ (specifically, two neurons). Here we consider that the link weight between two neurons is not only related to the neurons themselves, but also to the synapses between them. $\sigma(.)$ is the activate function.
$X_{i}=Y_{j}$ (16)
The link weights between neurons determine the extent of influence they have on each other. Larger link weights mean a tighter ”link” between two neurons. In everyday life, the brain continuously changes the link weights between neurons in response to inputs from the external world, allowing intelligence to emerge from the micro to the macro. Therefore, if researchers want to simulate the process of information propagation in the brain on large scale in a computer, determining the link weights between neurons is essential. However, existing experimental methods are not well-suited for directly measuring the weights between neurons in a whole brain, so adopting some alternative approach is necessary. Below, we introduce two hypotheses that use different methods to estimate the connection weights between neurons.
3.2.1 Average Hypothesis
$\omega(n_{k},e_{i},e_{j})=1$ (17)
Then the normalized equation is as follows.
$Y_{j}=\sigma(\frac{\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})}{{{Max}(\{\omega(n_{k},e_{i},e_{j})|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset,n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\})})$ (18)
3.2.2 Distance Hypothesis
The continuous evolution of organisms has led to the brain’s spatial structure being highly complex, with a small area possibly containing a large number of neurons. Besides, a fact is messages are transmitted through bioelectric currents. Assuming that the distance is longer, the resistance will be greater, resulting in a smaller communication effect. Based on this, there emerges a hypothesis: Is the strength of interactions between neurons inversely proportional to the spatial distance between them? So in the distance hypothesis we assume that the impact of input synapses on neurons is inversely proportional to the distance between them. Then the norm can be formulated as follows.
$\omega(n_{k},e_{i},e_{j})=(\parallel\vec{r_{k}}\parallel^{2}-\parallel\vec{r_{j}}\parallel^{2})^{\frac{1}{2}}$ (19)
Where ${r_{k}}$, ${r_{j}}$ are the center position vector of synapse and cell body, respectively.
Then the transmission function is defined as the following manner.
$Y_{j}=\sigma(\frac{\sum_{\{i|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset\}}\sum_{n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\sigma(\omega(n_{k},e_{i},e_{j})X_{k})}{{{Max}(\{\omega(n_{k},e_{i},e_{j})|e_{i}^{+}\cap e_{j}^{-}\neq\emptyset,n_{k}\in e_{i}^{+}\cap e_{j}^{-}}\})})$ (20)
3.3 Evaluation Criteria
In previous study, researchers used classic indicators such as signaling cost and activation time to evaluate the quality of the model[11][48][49]. However, these indicators can’t prove the model accords with the true communication process even if the activation time is short and the signaling cost is minimum, etc. To better assess whether a model can reflect the real situation of brain communication, we use the percentage of activated neurons in regions that should work under a certain stimulus. To illustrate, if the brain gets a visual stimulus, the region related to vision should work while the region not related to vision should keep in a resting state. Then we record the percentage of neurons activated and if regions related to the certain function have a significant higher activation percentage than the average activation percentage of the whole brain, we call it a quasi-real activation pattern.
We define the consistency indicator $C$. The consistency indicator $C$ in region $r$ which has a certain function (visual, olfactory, etc) is defined in the following manner.
$C_{r}=N_{r}^{act}/N_{r}$ (21)
3.4 Large Spatial Network Visualization
The visualization software of experiment observation on large spatial networks is important for researchers[50][51]. However, there is still a lack of visualization system that can observe information propagation on large spatial networks in real-time, which poses difficulties in experiment. To observe directly the communication progress on this super large Drosophila connectome, we develop a large spatial network visualization software.
The drawing part mainly employs the Canvas provided by HTML5 and we use the Three.js based on WebGL technology to achieve spatial network visualization effects, while the user interface display is constructed by Vue.js, providing users with an intuitive and interactive interface to display and analyze data. Fig. 5 is an example of the 3D visualization system. This software will be made available to other researchers with the author’s permission.
(a) Side view 1
(b) Side view 2
(c) Side view 3
(d) Top view 1
(e) Top view 2
(f) Top view 3
(g) Front view 1
(h) Front view 2
(i) Front view 3
Figure 5: Visualization of Spatial Drosophila Brain Network
3.4.1 Front End Technical Details
3.4.2 Back End Technical Details
4 Experimental Results and Analysis
In this part we introduce the experimental results and analysis. Above all, we introduce the activation pattern experiment. We give different stimuli and record the activation pattern of different neurons. In addition, we give a unilateral stimulus (stimulus one hemisphere) and record the bilateral response of neurons. Finally, to prove the network structure matters in brain network communication, we calculate the average physical and network distance between different areas.
4.1 Experimental Settings
The basic methodology of the experiment is described as follows. In this study, a continuous stimulus signal (binary 0-1 signal) is applied to a subset of input neurons (such as visual or olfactory neurons) in the brain. Subsequently, we compute the signal’s propagation process within the brain and record the responses of both intermediate and output neurons. The evaluation criterion is determined by calculating the ratio of neurons that ought to be activated to those that are actually activated in a specific area. Given that the brain functions as a type of shallow neural network[52], we document the outcomes at an iteration step size of 5, by which point the majority of neurons have achieved a state of stability. Moreover, as the input remains constant and no randomization parameters have been introduced, the outcomes of each experiment are consistent, eliminating the need for repetition.
Next, we will introduce the background knowledge of the neurons involved in this paper. Here are introductions of these neurons.
For further information, please read [43].
4.2 Activation Patterns on Different Type of Neurons
@@
Primarily, to validate the efficacy of our model in modifying signal transmission between input and output neurons, we showcase the activation patterns of intermediate and output neurons throughout the information transmission process. This demonstration serves to substantiate that our model accurately reflects real-world scenarios, shaping a quasi-real activation pattern.
@@
The specific idea for the experiment is following : we stimulate different types of neurons (such as visual input neurons, olfactory neurons, etc.), and record the rate that neurons that are activated in different brain areas. If a specific stimulus causes a significantly greater response than the average in the brain area that it is theoretically supposed to activate, then it is considered that the model can generate a quasi-real activation pattern in the brain. We give 2 types of input neurons (visual,olfactory) stimulus display 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory).
To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks. Specifically, we calculate the average network distance from input neurons (visual/olfactory) to corresponding neurons (optic, visual_projection/ALLN, ALPN, ALON, ALLN), round it to the nearest integer to determine the average layer of that neuron, and then randomly swap the connections between nodes within each layer.
@@
Fig. 6 shows the activation patterns of different models. The different neuron types are on the horizontal ordinate and the activation percentage is on the vertical ordinate. The horizontal line represents the percentage of activated neurons in the whole brain. From Fig. 6 we can find that under the average hypothesis, visual stimulus cause a strong impact (significantly more than average) on visual neurons (optic, visual projection) while causing a slight impact (significantly less than average) on other type of neurons. Specifically, under visual stimulus, in response to visual stimuli, the optic (vision related area) exhibit a 10.26% higher activation rate compared to the average level, while no olfactory-related neurons are activated. In addition, olfactory stimulus cause a strong impact (significantly bigger than average level) on olfactory neurons (ALLN, ALIN, ALPN, ALON), exceeding the average level by 81.49%. And it cause a slight impact (significantly less than average) on other type of neurons. For example, it only activates 0.04% of optic and 0.79% of visual_projection.
Furthermore, the results are not as significant on the distance model. Some incorrect areas were activated: The activate rate of ALLN is only 7.08%, with a 5.16% difference from the average level (Fig. 6(c)). Besides, ALIN, ALPN and ALLN, which should not be activated, have activate rates of 12.50%, 0.15%, and 1.89%, respectively (Fig. 6(c)). These results indicate that whether distance hypothesis can describe the actual situation remains to be discussed, which is a refutation for previous research[53].
As for the shuffle model, the ALIN, ALPN, ALON and ALLN which should not be activated have activate rates of 70.00%, 54.13%, 61.25% and 62.31%, respectively (Fig. 6(e)). Although these activate rates do not meet the average level, they are significant activation rates that means many neruons are activated incorrectly. Moreover, optic and visual_projection are activated incorrectly, with activate rates of 55.11% and 42.52%, respectively (Fig. 6(f)). Besides, ALLN, ALPN, ALON, and ALLN which are supposed to exceed the average level are actually lower than the average level by 0.01%, 17.74%, 21.25% and 9.34%, respectively (Fig. 6(f)).
These results indicate that after shuffling the network structure, the original quasi-real activation patterns completely disappear. It means the real network structure matters in generating the quasi-real activation patterns. In addition, even just designing simple models, it is feasible to simulate the propagation of information in the brain at the level of neurons and generate a quasi-real activation pattern.
(a) visual stimulus, average model
(b) olfactory stimulus, average model
(c) visual stimulus, distance model
(d) olfactory stimulus, distance model
(e) visual stimulus, shuffle model
(f) olfactory stimulus, shuffle model
@@
Figure 6: Experiment of intrinsic neurons’ activation. $\sigma_{0}=0.8$. For shuffle model, we conduct 5 repeated experiments and take the average and standard deviation.
@@ 4.3 Distances between Input and Activation Areas
In the previous section, the quasi-real activation patterns on different type of neurons of the brain was presented. Through observation, this relationship was identified, and a quasi-real activation pattern was observed in the drosophila brain. There is an issue here: Is the quasi-real activation pattern determined by the brain network’s structure, the spatial structure of the brain, or the quality of the network communication model? Based on the puzzle, we calculated the average network distance and spatial distance between different areas of the brain network. The optic and visual_projection area are related to vision and ALLN, ALIN, ALON, ALPN, LHCENT and LHLN are related to olfactory. The results are presented in Tab. 2.
@@
This results reflect the importance of network structure in neural networks, that is, the ability to reproduce certain functions of the brain by considering only the network structure.

### v3

@@ 2.2 Visualization and Statistic Properties of Drosophila Connectome
These statistic results are a valuable supplement to previous findings[48] and are crucial for for improved understanding of the Drosophila brain connectcome structure.
2.3 Activation Patterns under Different Type of Stimuli
@@
Primarily, to validate the efficacy of our model in signal transmission between input and output neurons, we showcase the activation patterns of intermediate and output neurons throughout the information transmission process. This demonstration serves to substantiate that our model accurately reflects real-world scenarios, shaping a quasi-real activation pattern.
@@
The specific idea for the experiment is following: we stimulate different types of neurons (such as visual input neurons, olfactory neurons, etc.), and record the rate that neurons that are activated in different brain areas. If a specific stimulus causes a significantly greater response than the average in the brain area that it is theoretically supposed to activate, then it is considered that the model can generate a quasi-real activation pattern in the brain. We give 2 types of input neurons (visual, olfactory) stimulus and record the activation patterns of 6 type of neurons: optic (visual), visual_projection (visual), ALLN (olfactory), ALIN (olfactory), ALPN (olfactory) and ALON (olfactory). When we give neuron a stimulus, we set the state of neuron as 1.
To demonstrate the importance of network structure, we add a comparative model that randomly reconnects networks. Specifically, we perform degree-preserving rewiring on a certain proportion of edges in the brain, which means perturbing the network structure while ensuring that the degree of each node remains unchanged. In Fig.2 we set reconnect rate $p=10\%$. For its sensitivity analysis, please read appendix S4.
@@
Fig.2 shows the activation patterns of different models. From Fig. 2(a), (e), and (i), it shows visual stimulus cause a strong impact (significantly more than average) on visual neurons (optic, visual projection) while causing a slight impact (significantly less than average) on other type of neurons under all three kinds of models. Specifically, under visual stimulus, in response to visual stimuli, the optic (vision related area) exhibit a 36.39%, 4.69% and 9.69% ’s activate rates, respectively. At the mean time, no olfactory-related neurons are activated.
In addition, olfactory stimulus cause a strong impact (significantly bigger than average level) on olfactory neurons (ALLN, ALIN, ALPN, ALON), reaching the activation rate by 94.07%, 78.01%, and 69.75%, respectively. And it cause a slight impact (significantly less than average) on other type of neurons. For example, it only activates 0.04% of optic and 0.79% of visual_projection in threshold model (Fig.2(c)).
Although the activation rates of the three models in specific areas differ (due to different activation mechanisms), the significant difference in the activation ratio between correlated and uncorrelated regions indicates the emergence of similar quasi-real activation patterns across all models. This suggests that different activation mechanisms are not the fundamental reason for changes in network functionality.
As for the disturbing models, in Fig.2 (b), (d), (f), (h), (j) and (l), the quasi-real activation patterns totally disappear. Across all activation modes and all types of stimuli, no significant differences in activation regions were observed. For example, in the threshold model, the maximum difference in the activation ratios of all areas under visual stimuli was only 3.162%. Moreover, the range of the regional activation ratios versus the average activation ratio was only 6.32%. Additionally, in the sigmoid and LIF models, no significant activation was observed in any area, indicating that stimuli under network disturbance did not lead to any significant response in any area.
These results indicate that no matter neuron model is, it will lead to similar activation patterns which correspond to reality. After disturbing the network structure, the original quasi-real activation patterns in all models completely disappear. It means the real network structure matters than activation mechanisms in generating the quasi-real activation patterns. Different activation mechanisms are not the primary cause of this realistic activation mechanism; it is the brain’s unique network structure that plays the crucial role. This aligns with the studies of complex systems: individual capabilities are not the determining factors in complex systems; what matters is the organization and interaction structure among individuals. This perspective highlights the importance of individual relationships and the overall system structure in determining emergent phenomena.
(a) threhold, visual
(b) threhold, visual (after disturb)
(c) threhold, olfactory
(d) threhold, olfactory (after disturb)
(e) sigmoid, visual
(f) sigmoid, visual (after disturb)
(g) sigmoid, olfactory
(h) sigmoid, olfactory (after disturb)
(i) LIF, visual
(j) LIF, visual (after disturb)
(k) LIF, olfactory
(l) LIF, olfactory (after disturb)
@@
Figure 2: Experiment of intrinsic neurons’ activation. For disturbing model, we conduct 5 repeated experiments and take the average and standard deviation.
@@ 2.4 Distances between Input and Activation Areas
To further investigate how the network structure, rather than the spatial distribution of neurons, or the parameters of the network model or activation mechanisms, predominantly determines the dynamics of brain activity, we calculated the average network distance and spatial distance between different areas of the brain network. The optic and visual_projection area are related to vision and ALLN, ALIN, ALON, ALPN, LHCENT and LHLN are related to olfactory. The results are presented in Tab. 2.
@@
This results reflect the importance of network structure in neural networks, that is, the ability to reproduce certain functions of the brain by considering the network structure only.